# Gradient elasticity in Swift-Hohenberg and phase-field crystal models

### v1

@@ Abstract
The Swift-Hohenberg (SH) and Phase-Field Crystal (PFC) models are minimal yet powerful approaches for studying phenomena such as pattern formation, collective order, and defects via smooth order parameters. They are based on a free-energy functional that includes elasticity with linear, nonlinear, and strain-gradient contributions. The latter is peculiar to gradient elasticity (GE), a theory that accounts for elasticity effects at microscopic scales by introducing additional characteristic lengths. While numerical simulations of SH and PFC models display some distinctive GE features, such as the regularization of stress singularities at crystalline defects, a comprehensive analysis has yet to be conducted. This study addresses how GE is incorporated into SH and PFC models. We calculate its characteristic lengths for various lattice symmetries in an approximated setting. We then discuss the effective elasticity theory encoded in SH and PFC models via numerical simulations of dislocation stress fields and comparisons with solutions within first and second strain GE. We then demonstrate that effective GE characteristic lengths depend on the free-energy parameters, similar to the correlation length, thus unveiling how they change with the quenching depth and the phenomenological temperature entering the considered models. The findings presented in this study enable a thorough discussion and analysis of small-scale elasticity effects in pattern formation using SH and PFC models and complete the elasticity analysis therein. Additionally, we provide a microscopic foundation for GE in the context of order-disorder phase transitions.
@@ 1 Introduction
Gradient elasticity (GE) is a theoretical framework that extends classical continuum mechanics with additional higher-order derivatives to capture the effects of small length scales[1, 2, 3, 4, 5]. It incorporates additional characteristic lengths accounting for internal material structure and discreteness, thereby including size effects relevant at the nano- and micro-scale. Such a theory has been tested and exploited in different frameworks, e.g., classical mechanics and dislocation dynamics, and applied in contexts such as flexoelectricity and biomechanics [6, 7, 8, 9, 10, 11, 12]. It has also been extended to describe strain gradient plasticity [13, 14, 15, 16]. Parameters entering the theory have been connected to interatomic potentials and, in general, can be determined by atomistic methods [17, 18, 19]. A key feature of GE is that it overcomes some limitations of continuum elasticity theory, such as unphysical singularities emerging from approximated descriptions of small scales. This leads, for instance, to regularized elastic fields at the dislocation core [20, 21, 22, 23], which are instead singular in classic linear elasticity [24].
@@
Elasticity-driven pattern formation emerges in physical systems due to the interplay between mechanical deformation and other physical contributions such as capillarity, solidification, crystal growth dynamics, and grand potential jumps across interfaces. This can occur in various systems such as hard and soft materials as well as biological tissues [25, 26]. Concerning the study of general features of pattern formation, a central role has been played by the so-called Swift-Hohenberg (SH) model [27, 28] and its extensions. It is based on a partial differential equation (the SH equation) that describes the spatiotemporal evolution of a real order parameter field, $\psi$, comprising linear differential operators and nonlinear polynomials. Such an equation favors the linear growth of periodic modes of $\psi$. At the same time, nonlinear terms lead to saturation of this growth and selection of patterns with different symmetry depending on parameters. Notably, the SH equation can be written in a variational form through the definition of a free energy (or Ljapunov) functional $F[\psi]$ as $\partial_{t}\psi=-\delta F/\delta\psi$. Therefore, stable stationary states correspond to the minima of such functional. While it was initially proposed to model the so-called Rayleigh-Benard convection in a heated fluid, similar equations have been exploited to describe the emergence of patterns involving elasticity effects in different contexts [29, 30, 31]. The connection of SH-like models to elasticity became evident in the conservative formulation, the so-called phase field crystal (PFC) model, explicitly introduced to model elasticity in crystal growth [32, 33, 34, 35]. In its basic formulation, the PFC model is based on a free energy analogous to the SH model, albeit with conservative dynamics $\partial_{t}\psi=\nabla^{2}(\delta F/\delta\psi)$, where $\psi$ can be considered as the time average of microscopic atomic density over vibrational timescales. The periodic patterns are then representative of lattices in crystalline materials. Such a model has been successfully applied to describe mesoscale phenomena in crystals, including elasticity effects during crystal growth, dislocation dynamics, and microstructure evolution [33, 35, 36, 37, 38] also with non-conservative formulations which, de facto, correspond to the SH equation [39].
@@
Elasticity is thus naturally encoded in SH/PFC models. By small perturbation of the periodicity of the order parameter, the energy can be written as a quadratic form of the strain [33, 40] resembling classical models of linear elasticity. Elastic constants are then related to the lattice symmetry described by $\psi$, and their magnitudes depend on the model parameters. Advanced models have also been proposed to account for elastic relaxation timescales [41, 40, 42, 43, 44]. Without approximations, the elastic energy that can be written upon displacing the density field also comprises higher-order terms, including nonlinear and strain-gradient terms. While the former have been the object of dedicated analyses[45, 46, 47], the latter have typically been neglected. However, numerical simulations exhibit features typical of GE. For instance, the stress field at dislocations in PFC simulations matches classical continuum elasticity in the far field but is nonsingular at their cores [48, 49, 50, 47]. This is expected from a qualitative point of view as the fields $\psi$ are smooth functions of spatial coordinates due to the underlying free energy functionals. However, the emerging regularization explicitly resembles that obtained in GE theories [20], for which preliminary evidence has been reported [50, 47]. In this work, we study how GE is encoded in SH and PFC models. The importance of establishing this connection is two-fold. On the one hand, related effects are present in numerical simulations. A comprehensive understanding of elasticity in the considered models allows for a better interpretation of simulation results and enables more rigorous applications to small-scale elasticity problems. On the other hand, the SH and PFC models provide microscopic theories for systems exhibiting order at a small scale, like reaction-diffusion systems and crystalline materials. As shown below, GE emerges naturally from the perturbation of their ground states. Therefore, GE is here self-consistently obtained as an elasticity theory valid at a small scale and connected to a prototypical, general framework for order-disorder phase transition.
@@
This paper is organized as follows. In Sect. 2, we recall the basics of the GE theory, reporting the key information exploited in the current analysis. In Sect. 3, we first illustrate the SH free energy functional. Then we show how, in a simplified setting, GE naturally emerges from it upon perturbation of the microscopic density for both stripe (Sect. 4.1) and crystalline (Sect. 4.2) phases, also discussing extended parameterizations via the direct definition of a two-point correlation function (Sect. 4.3). In Sect. 5, we address in detail the effective GE theory that transpires without approximations in numerical simulations, focusing on the relevant case of the elastic field at dislocations. Therein, we also discuss the dependence of the characteristic lengths on model parameters and their implications. Conclusions are summarized in Sect. 6.
@@ 2 Gradient Elasticity
In first strain-gradient elasticity (GE-1), the elastic energy density accounts for strain and its first derivatives. In Toupin–Mindlin theory of anisotropic GE-1 [1, 2], it generally reads 111we hereafter adopt Einstein summation convention
@@
with $\bm{\varepsilon}=\frac{1}{2}(\nabla\mathbf{u}+\nabla\mathbf{u}^{\intercal})$ the strain tensor, $\mathbf{u}$ the macroscopic displacement field, and $\mathbb{C}$, $\mathbb{E}$, and $\mathbb{D}$, the constitutive four-, five- and six-rank tensors encoding elastic constants. This formulation simplifies the more general Mindlin’s theory of elasticity in the presence of microstructures, which includes quantities at both macro and microscale [3, 5]. This multiscale formulation, however, can hardly be exploited for practical purposes, and gradient elasticity theories based on the macroscopic displacement $\mathbf{u}$ are usually considered. In practice, the role played by microscopic deformation gradients is encoded by second gradients of the macroscopic displacement (Form I) or first gradients of the macroscopic strain (Form II), the latter leading to Eq. (1). $\mathbb{C}$ and $\mathbb{D}$ possess both minor and major symmetries, while $\mathbb{E}$ has only minor symmetries. Illustrated for $\mathbb{C}$, these symmetries read component-wise
@@
$\tau_{ijm}$ $=\frac{\partial w}{\partial(\partial_{m}\varepsilon_{ij})}=\mathbb{E}_{klijm}(\partial_{m}\varepsilon_{kl})+\mathbb{D}_{ijmkln}(\partial_{n}\varepsilon_{kl}).$ (5)
@@
which can be recast in a field equation for the displacement:
@@
In the following, we will also use the Voigt notation for the fourth-order tensor $\mathbb{C}$ using
@@
and for the six-order tensor $\mathbb{D}$ using
@@ 2.1 Isotropic materials
For isotropic materials,
$\mathbb{C}_{ijkl}=\lambda\delta_{ij}\delta_{kl}+\mu(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}),$ (11)
@@
$\begin{split}\mathbb{D}_{ijmkln}&=\frac{a_{1}}{2}(\delta_{ij}\delta_{km}\delta_{ln}+\delta_{ij}\delta_{kn}\delta_{lm}+\delta_{kl}\delta_{im}\delta_{jn}+\delta_{kl}\delta_{in}\delta_{jm})+2a_{2}\delta_{ij}\delta_{kl}\delta_{mn}\\ &+\frac{a_{3}}{2}(\delta_{jk}\delta_{im}\delta_{ln}+\delta_{ik}\delta_{jm}\delta_{ln}+\delta_{il}\delta_{jm}\delta_{kn}+\delta_{jl}\delta_{im}\delta_{kn})+a_{4}(\delta_{il}\delta_{jk}\delta_{mn}+\delta_{ik}\delta_{jl}\delta_{mn})\\ &+\frac{a_{5}}{2}(\delta_{jk}\delta_{in}\delta_{lm}+\delta_{ik}\delta_{jn}\delta_{lm}+\delta_{jl}\delta_{km}\delta_{in}+\delta_{il}\delta_{km}\delta_{jn}),\end{split}$ (12)
@@
and
$L_{ik}^{\mathbb{C}}$ $=(\lambda+2\mu)\partial_{i}\partial_{k}+\mu(\delta_{ik}-\partial_{i}\partial_{k}),$ (14)
$L_{ik}^{\mathbb{D}}$ $=2(a_{1}+a_{2}+a_{3}+a_{4}+a_{5})\partial_{i}\partial_{k}+\frac{1}{2}(a_{3}+2a_{4}+a_{5})(\delta_{ik}-\partial_{i}\partial_{k}).$ (15)
One may redefine $L_{ik}$ as
$L_{ik}=(\lambda+2\mu)[1-\ell_{1}^{2}\nabla^{2}]\partial_{i}\partial_{k}+\mu[1-\ell^{2}_{2}\nabla^{2}](\delta_{ik}\nabla^{2}-\partial_{i}\partial_{k}),$
@@ 2.2 Centrosymmetric materials
For centrosymmetric materials,
$\mathbb{C}_{ijkl}=C_{12}\delta_{ij}\delta_{kl}+C_{44}(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})+(C_{11}-C_{12}-2C_{44})\delta_{ijkl},$ (18)
@@
The constants entering the prefactor of $\delta_{ijkl}$ are typically used to quantify the anisotropy of a cubic material alternatively, e.g., via the so-called Zener ratio $z=2C_{44}/(C_{11}-C_{12})$ with $z=1$ for isotropic materials, for which Eq. (18) reduces to Eq. (11). The strain-gradient elasticity tensor can be written as
$\begin{split}\mathbb{D}_{ijmkln}&=\frac{a_{1}}{2}(\delta_{ij}\delta_{km}\delta_{ln}+\delta_{ij}\delta_{kn}\delta_{lm}+\delta_{kl}\delta_{im}\delta_{jn}+\delta_{kl}\delta_{in}\delta_{jm})+2a_{2}\delta_{ij}\delta_{kl}\delta_{mn}\\ &+\frac{a_{3}}{2}(\delta_{jk}\delta_{im}\delta_{ln}+\delta_{ik}\delta_{jm}\delta_{ln}+\delta_{il}\delta_{jm}\delta_{kn}+\delta_{jl}\delta_{im}\delta_{kn})+a_{4}(\delta_{il}\delta_{jk}\delta_{mn}+\delta_{ik}\delta_{jl}\delta_{mn})\\ &+\frac{a_{5}}{2}(\delta_{jk}\delta_{in}\delta_{lm}+\delta_{ik}\delta_{jn}\delta_{lm}+\delta_{jl}\delta_{km}\delta_{in}+\delta_{il}\delta_{km}\delta_{jn})\\ &+a_{6}(\delta_{ik}\delta_{jlmn}+\delta_{il}\delta_{jkmn}+\delta_{jk}\delta_{ilmn}+\delta_{jl}\delta_{ikmn})\\ &+a_{7}(\delta_{km}\delta_{ijln}+\delta_{lm}\delta_{ijkn}+\delta_{in}\delta_{jklm}+\delta_{jn}\delta_{iklm})\\ &+a_{8}\delta_{mn}\delta_{ijkl}+a_{9}(\delta_{ij}\delta_{klmn}+\delta_{kl}\delta_{ijmn})\\ &+a_{10}(\delta_{im}\delta_{jkln}+\delta_{jm}\delta_{ikln}+\delta_{kn}\delta_{ijlm}+\delta_{ln}\delta_{ijkm})+a_{11}\delta_{ijklmn}\end{split}$ (20)
Similarly to the case of isotropic material, one may define
@@
entering the operator $L_{ik}$ for the generalized equations for displacements [21]:
$\begin{split}L_{ik}&=(C_{11}+2C_{44})[1-\ell_{1}^{2}\nabla^{2}]\partial_{i}\partial_{k}+C_{44}[1-\ell^{2}_{2}\nabla^{2}](\delta_{ik}\nabla^{2}-\partial_{i}\partial_{k})\\ &+(C_{11}-C_{12}-2C_{44})[1-\ell^{2}_{3}\nabla^{2}]\delta_{ik}\partial_{i}\partial_{k}-(a_{6}\delta_{ik}\delta_{jlmn}+a_{11}\delta_{ijklmn})\partial_{j}\partial_{l}\partial_{m}\partial_{n}\\ &-(a_{6}+a_{7}+a_{9}+a_{10})(\delta_{klmn}\partial_{i}+\delta_{ilmn}\partial_{k})\partial_{l}\partial_{m}\partial_{n}.\end{split}$ (22)
In this case, terms explicitly depending on $a_{i}$ appear in $L_{ik}$, marking the inherent anisotropy of cubic crystals.
2.3 Special second-strain gradient elasticity
In second strain-gradient elasticity (GE-2), second derivatives of the strain field are retained in the elastic energy density, $w\equiv w(\varepsilon_{ij},\partial_{k}\varepsilon_{ij},\partial_{k}\partial_{\ell}\varepsilon_{ij})$. For isotropic materials, a special formulation can be considered still encoding two characteristic lengths $\varrho_{1,2}$. They enter the mechanical equilibrium condition as [21]

### v2

@@ Abstract
The Swift-Hohenberg (SH) and Phase-Field Crystal (PFC) models are minimal yet powerful approaches for studying phenomena such as pattern formation, collective order, and defects via smooth order parameters. They are based on a free-energy functional that inherently includes elasticity effects. This study addresses how gradient elasticity (GE), a theory that accounts for elasticity effects at microscopic scales by introducing additional characteristic lengths, is incorporated into SH and PFC models. After presenting the fundamentals of these theories and models, we first calculate the characteristic lengths for various lattice symmetries in an approximated setting. We then discuss numerical simulations of stress fields at dislocations and comparisons with analytic solutions within first and second strain-gradient elasticity. Effective GE characteristic lengths for the elastic fields induced by dislocations are found to depend on the free-energy parameters in the same manner as the phase correlation length, thus unveiling how they change with the quenching depth. The findings presented in this study enable a thorough discussion and analysis of small-scale elasticity effects in pattern formation and crystalline systems using SH and PFC models and, importantly, complete the elasticity analysis therein. Additionally, we provide a microscopic foundation for GE in the context of order-disorder phase transitions.
@@ 1 Introduction
Gradient elasticity (GE) is a theoretical framework that extends classical continuum mechanics with additional higher-order derivatives to capture the effects of small length scales [1, 2, 3, 4, 5, 6, 7]. It incorporates additional characteristic lengths accounting for internal material structure and discreteness, thereby allowing the description of size effects relevant at the nano- and micro-scale. Such a theory has been tested and exploited in different frameworks, e.g., classical mechanics and dislocation dynamics, and applied in contexts such as flexoelectricity and biomechanics [8, 9, 10, 11, 12, 13, 14]. It has also been extended to describe strain-gradient plasticity [15, 16, 17, 18]. Parameters entering the theory have been connected to interatomic potentials and, in general, can be determined by atomistic methods [19, 20, 21]. A key feature of GE is that it overcomes some limitations of continuum elasticity theory, such as nonphysical singularities emerging from approximated descriptions of small scales. This leads, for instance, to regularized elastic fields at the dislocation core [22, 23, 24, 25, 26], which are instead singular in classic linear elasticity [27].
@@
Elasticity-driven pattern formation emerges in a variety of physical systems—e.g., hard and soft materials, biological tissues [28, 29]—due to the interplay between mechanical deformation and other physical contributions such as capillarity, solidification, crystal growth dynamics, and grand potential jumps across interfaces. In the study of general features of pattern formation, a central role has been played by the so-called Swift-Hohenberg (SH) model [30, 31] and its extensions. It describes the spatiotemporal evolution of a real order parameter field $\psi$ based on a partial differential equation comprising linear differential operators and polynomial terms (the SH equation). Such an equation favors the linear growth of periodic modes of $\psi$ until the nonlinear contributions induce saturation, leading to the selection of patterns with different symmetries depending on the parametrization. Notably, the SH equation can be written in a variational form through the definition of a free energy (or Lyapunov) functional $F[\psi]$ as $\partial_{t}\psi=-\delta F/\delta\psi$. Therefore, stable stationary states correspond to the minima of such functional. While it was initially proposed to model the so-called Rayleigh-Bénard convection in a heated fluid, similar equations have been exploited to describe the emergence of patterns involving elasticity effects in different contexts [32, 33, 34]. The connection of SH-like models to elasticity became evident in the conservative formulation, the so-called phase-field crystal (PFC) model, explicitly introduced to model elasticity in crystal growth [35, 36, 37, 38]. In its basic formulation, the PFC model is based on a free energy analogous to the SH model, albeit with conservative dynamics $\partial_{t}\psi=\nabla^{2}(\delta F/\delta\psi)$, where $\psi$ can be considered as the time average of microscopic atomic density over vibrational timescales. The periodic pattern, in that case, represents the lattice structure of crystalline materials. Such a model has been successfully applied to describe mesoscale phenomena in crystals, including elasticity effects during crystal growth, dislocation dynamics, and microstructure evolution [36, 38, 39, 40, 41].
@@
Elasticity is thus naturally encoded in SH/PFC models. Introducing a small perturbation in the periodicity of the order parameter and neglecting higher-order terms, the energy can be written as a quadratic form of the strain [36, 42] resembling classical models of linear elasticity. Elastic constants are then related to the lattice symmetry described by $\psi$, with magnitudes dependent on the model parameters. Advanced models have also been proposed to account for elastic relaxation timescales [43, 42, 44, 45, 46]. Without approximations, the elastic energy obtained by perturbing the density field also comprises higher-order terms, including nonlinear and strain-gradient terms. While nonlinearities have been the object of dedicated analyses [47, 48, 49], strain-gradient terms have typically been neglected. However, numerical simulations exhibit features typical of GE. For instance, the stress field at dislocations in PFC simulations matches classical continuum elasticity in the far field but is nonsingular at the core[50, 51, 52, 49]. Qualitatively, some form of regularization is indeed expected as the fields $\psi$ are smooth functions of the spatial coordinates due to the underlying free energy functionals. However, the emerging regularization explicitly resembles that obtained in GE theories [22], for which preliminary evidence has been reported [52, 49].
In this work, we study how GE is encoded in SH/PFC models. We bridge for the first time the field of nonlocal elasticity encoding microscopic effects (as described by GE) with the theory of pattern formation at the micro-to-meso-scale conveyed by SH/PFC models, which also serve as convenient frameworks for numerical simulations. The importance of establishing this connection is thus two-fold. (i) Current approaches in the field of GE typically assume some form of elastic energy; here, we show how GE free energies and constants entering this theory follow from microscopic models like SH and PFC, as well as related aspects, like the regularization of elastic fields at defects. (ii) By showing analytically and numerically that strain-gradient terms are inherently encoded in the SH free energy, we generally assess that approaches based on SH and PFC models retain these effects and discuss how GE emerges with concrete examples. Our work then enables advanced interpretation of simulation results. It paves the way for using SH and PFC models to study GE effects self-consistently, i.e., following assumptions on microscopic ordering in lattices or, more generally, patterns retaining some degree of order.
@@
This paper is organized as follows. In Sect. 2, we recall the basics of the GE theory, reporting the key information exploited in the current analysis. Readers familiar with this theory may skip this section and use it as a reference for the adopted notation. Similarly, in Sect. 3, we illustrate the SH free energy functional and recall the basics of SH and PFC models. From Sect. 4, we discuss novel aspects and results. As first main focus, we discuss how GE naturally arises from the SH energy functional upon perturbation of the microscopic density (Sect. 4.1). We then analyze in detail the derived theory for stripe (Sect. 4.2) and crystalline (Sect. 4.3) phases, followed by a discussion of the results and their implications (Sect. 4.4). While this analysis is obtained for a minimal formulation of the SH energy functional, extended parametrizations achieved via the definition of a two-point correlation function are also discussed (Sect. 4.5). As the second main focus, we address in Sect. 5 the effective GE theory that transpires without approximations in numerical simulations. We study, in particular, the relevant case of the elastic fields at dislocations. After reporting the details concerning numerical simulations (Sect. 5.1), we characterize the stress field and its regularization at the core of an edge dislocation via comparisons with analytic solutions from first and second strain-gradient elasticity and fitting of effective characteristic lengths (Sect. 5.2). Results including the dependence of the characteristic lengths on model parameters, like the quenching depth, and their implications, are then discussed (Sect. 5.3). Conclusions are summarized in Sect. 6. Technical details concerning numerical methods, analytic expressions for the stress fields of dislocations in first and second strain-gradient elasticity, details of fitting procedures, and additional information, are reported in the Appendix.
@@ 2 Gradient Elasticity
In first strain-gradient elasticity (GE-1), the elastic energy density accounts for strain and its first derivatives. In the Toupin–Mindlin theory of anisotropic GE-1 [1, 2], following the convenient notation of Refs. [6, 7], the energy density generally reads  [^1] We hereafter adopt Einstein summation convention.
@@
with $\bm{\varepsilon}=\frac{1}{2}(\nabla\mathbf{u}+\nabla\mathbf{u}^{\intercal})$ the strain tensor, $\mathbf{u}$ the macroscopic displacement field, and $\mathbb{C}$, $\mathbb{E}$, and $\mathbb{D}$, the constitutive four-, five- and six-rank tensors encoding elastic constants. This formulation simplifies the more general Mindlin’s theory of elasticity in the presence of microstructures, which includes quantities at both macro- and micro-scale [3, 5]. This multiscale formulation, however, can hardly be exploited for practical purposes, and gradient elasticity theories based on the macroscopic displacement $\mathbf{u}$ are usually considered. In practice, the role played by microscopic deformation gradients is encoded by second gradients of the macroscopic displacement (Form I) or first gradients of the macroscopic strain (Form II), the latter leading to Eq. (1). $\mathbb{C}$ and $\mathbb{D}$ possess both minor and major symmetries, while $\mathbb{E}$ has only minor symmetries. Illustrated for $\mathbb{C}$, these symmetries read component-wise
@@
$\tau_{ijm}$ $=\frac{\partial w}{\partial(\partial_{m}\varepsilon_{ij})}=\mathbb{E}_{klijm}\varepsilon_{kl}+\mathbb{D}_{ijmkln}(\partial_{n}\varepsilon_{kl}).$ (5)
@@
which can be recast in a field equation for the displacement [6]:
@@
In the following, Voigt notation is also used for the fourth-order tensor $\mathbb{C}$
@@
with components denoted $C_{ij}$, and for the six-order tensor $\mathbb{D}$
@@ 2.1 Isotropic materials
For isotropic materials (iso), the elastic tensors read
$\mathbb{C}_{ijkl}^{\rm iso}=\lambda\delta_{ij}\delta_{kl}+\mu(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}),$ (11)
@@
$\begin{split}\mathbb{D}_{ijmkln}^{\rm iso}&=\frac{a_{1}}{2}(\delta_{ij}\delta_{km}\delta_{ln}+\delta_{ij}\delta_{kn}\delta_{lm}+\delta_{kl}\delta_{im}\delta_{jn}+\delta_{kl}\delta_{in}\delta_{jm})+2a_{2}\delta_{ij}\delta_{kl}\delta_{mn}\\ &+\frac{a_{3}}{2}(\delta_{jk}\delta_{im}\delta_{ln}+\delta_{ik}\delta_{jm}\delta_{ln}+\delta_{il}\delta_{jm}\delta_{kn}+\delta_{jl}\delta_{im}\delta_{kn})+a_{4}(\delta_{il}\delta_{jk}\delta_{mn}+\delta_{ik}\delta_{jl}\delta_{mn})\\ &+\frac{a_{5}}{2}(\delta_{jk}\delta_{in}\delta_{lm}+\delta_{ik}\delta_{jn}\delta_{lm}+\delta_{jl}\delta_{km}\delta_{in}+\delta_{il}\delta_{km}\delta_{jn}),\end{split}$ (12)
@@
and the operators
$L_{ik}^{\mathbb{C},\,{\rm iso}}$ $=(\lambda+2\mu)\partial_{i}\partial_{k}+\mu(\delta_{ik}\Delta-\partial_{i}\partial_{k}),$ (14)
$L_{ik}^{\mathbb{D},\,{\rm iso}}$ $=2(a_{1}+a_{2}+a_{3}+a_{4}+a_{5})\partial_{i}\partial_{k}\Delta+\frac{1}{2}(a_{3}+2a_{4}+a_{5})(\delta_{ik}\Delta-\partial_{i}\partial_{k}).$ (15)
Equations (14) and (15) correspond to the expressions in (8) with (isotropic) elastic constants as in (11) [6, 7]. Moreover, one may redefine $L_{ik}^{\rm iso}$ as
$L_{ik}^{\rm iso}=(\lambda+2\mu)[1-\ell_{1}^{2}\nabla^{2}]\partial_{i}\partial_{k}+\mu[1-\ell^{2}_{2}\nabla^{2}](\delta_{ik}\nabla^{2}-\partial_{i}\partial_{k}),$
@@
$\ell_{i}$ can be considered measures of the material’s internal characteristic lengths, such as the distance over which microstructural effects (like atomic lattice or grain structure) significantly influence the material’s response to deformation. In essence, they provide a way to include the effect of the material’s microstructure in continuum-scale models. $a_{i}$ may be expressed in terms of the components of $\mathbb{D}$, and computed by atomistic methods [7]. In the following, they are directly extracted by comparison of the considered elastic energy densities with the form given in (13), see also Sect. 4.1.
@@ 2.2 Centrosymmetric materials
Centrosymmetric materials (cs) possess a lattice structure with an inversion center $\mathbf{r}_{0}$, i.e., points $\mathbf{r}_{0}\pm\mathbf{r}$ are equivalent. For these materials, elastic constants read
$\mathbb{C}_{ijkl}^{\rm cs}=C_{12}\delta_{ij}\delta_{kl}+C_{44}(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})+(C_{11}-C_{12}-2C_{44})\delta_{ijkl},$ (18)
@@
The constants entering the prefactor of $\delta_{ijkl}$ are typically used to quantify the anisotropy of a cubic material via, e.g., the so-called Zener ratio $z=2C_{44}/(C_{11}-C_{12})$. For isotropic materials, $z=1$ and Eq. (18) reduces to Eq. (11) as expected. The strain-gradient elasticity tensor can be written as [7]
$\begin{split}\mathbb{D}_{ijmkln}^{\rm cs}&=\frac{a_{1}}{2}(\delta_{ij}\delta_{km}\delta_{ln}+\delta_{ij}\delta_{kn}\delta_{lm}+\delta_{kl}\delta_{im}\delta_{jn}+\delta_{kl}\delta_{in}\delta_{jm})+2a_{2}\delta_{ij}\delta_{kl}\delta_{mn}\\ &+\frac{a_{3}}{2}(\delta_{jk}\delta_{im}\delta_{ln}+\delta_{ik}\delta_{jm}\delta_{ln}+\delta_{il}\delta_{jm}\delta_{kn}+\delta_{jl}\delta_{im}\delta_{kn})+a_{4}(\delta_{il}\delta_{jk}\delta_{mn}+\delta_{ik}\delta_{jl}\delta_{mn})\\ &+\frac{a_{5}}{2}(\delta_{jk}\delta_{in}\delta_{lm}+\delta_{ik}\delta_{jn}\delta_{lm}+\delta_{jl}\delta_{km}\delta_{in}+\delta_{il}\delta_{km}\delta_{jn})\\ &+a_{6}(\delta_{ik}\delta_{jlmn}+\delta_{il}\delta_{jkmn}+\delta_{jk}\delta_{ilmn}+\delta_{jl}\delta_{ikmn})\\ &+a_{7}(\delta_{km}\delta_{ijln}+\delta_{lm}\delta_{ijkn}+\delta_{in}\delta_{jklm}+\delta_{jn}\delta_{iklm})\\ &+a_{8}\delta_{mn}\delta_{ijkl}+a_{9}(\delta_{ij}\delta_{klmn}+\delta_{kl}\delta_{ijmn})\\ &+a_{10}(\delta_{im}\delta_{jkln}+\delta_{jm}\delta_{ikln}+\delta_{kn}\delta_{ijlm}+\delta_{ln}\delta_{ijkm})+a_{11}\delta_{ijklmn}.\end{split}$ (20)
With these elasic constants, the operator $L_{ik}^{\rm cs}$ for the generalized equations for displacements (8) becomes [7]:
$\begin{split}L_{ik}^{\rm cs}&=(C_{11}+2C_{44})[1-\ell_{1}^{2}\nabla^{2}]\partial_{i}\partial_{k}+C_{44}[1-\ell^{2}_{2}\nabla^{2}](\delta_{ik}\nabla^{2}-\partial_{i}\partial_{k})\\ &+(C_{11}-C_{12}-2C_{44})[1-\ell^{2}_{3}\nabla^{2}]\delta_{ik}\partial_{i}\partial_{k}-(a_{6}\delta_{ik}\delta_{jlmn}+a_{11}\delta_{ijklmn})\partial_{j}\partial_{l}\partial_{m}\partial_{n}\\ &-(a_{6}+a_{7}+a_{9}+a_{10})(\delta_{klmn}\partial_{i}+\delta_{ilmn}\partial_{k})\partial_{l}\partial_{m}\partial_{n}.\end{split}$ (21)
with characteristic lengths
@@
In this case, terms explicitly depending on $a_{i}$ appear in $L_{ik}^{\rm cs}$, marking the intrinsic anisotropy of cubic crystals.
2.3 Special second strain-gradient elasticity
In second strain-gradient elasticity (GE-2), second derivatives of the strain field are retained in the elastic energy density, $w\equiv w(\varepsilon_{ij},\partial_{k}\varepsilon_{ij},\partial_{k}\partial_{\ell}\varepsilon_{ij})$. For isotropic materials, a formulation relying on two characteristic lengths $\varrho_{1,2}$ may be considered. They enter the mechanical equilibrium condition as [23]